a straight line with the slope 5 contains the points (1,2) and (3,K). Find the value of K.
Answer by xcentaur(357)

slope m of the line = difference in y-coordinates/ difference in x-coordinates


m = k-2/3-1
given that slope is 5,
k-2/3-1=5
k-2/2=5
k-2=10
k=12


Hope this helps,
Prabhat